OpenStudy (anonymous):

determine the shortest distance from the point (6,3) to the line y=-2x+1

OpenStudy (anonymous):

the shortest distance is the perpendicular distance which is 14/sqrt(5)

OpenStudy (anonymous):

two ways, calc way and thinking way

OpenStudy (anonymous):

basically we shld find the right bisector rite?

OpenStudy (anonymous):

thinking way is to realize that the line connecting the point and the other line giving the shortest distance will be perpendicular to it and therefore have slope of \[\frac{1}{2}\]

OpenStudy (anonymous):

so you have a line with slope \[\frac{1}{2}\] through (6,3). find the equation of the line and then see where they intersect by setting the two lines equal

OpenStudy (anonymous):

the equation is \[y-3=\frac{1}{2}(x-6)\] \[y-3=\frac{1}{2}x-3\] \[y=\frac{1}{2}x\]

OpenStudy (anonymous):

now solve \[-2x+1=\frac{1}{2}x\]

OpenStudy (anonymous):

\[x=\frac{2}{5}\]

OpenStudy (anonymous):

so \[y=\frac{1}{5}\]

OpenStudy (anonymous):

then use distance formula.

OpenStudy (anonymous):

thank u